Solve the differential equation.
(D²+D+D')z=0
1
Expert's answer
2020-08-06T16:06:33-0400

Let,"D=h,D'=k" ,thus

"(D²+D+D')z=0\cong h^2+h+k=0"

Hence,

"h_1=\frac{-1+\sqrt{1-4k}}{2}\\h_2=\frac{-1-\sqrt{1-4k}}{2}"

Then

"C.F=\sum C_1e^{h_1x+ky}+\sum C_2e^{h_2x+ky}"